Circular No. 8302

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ar, 2004as, AND 2004at A. Connolly, University of Pittsburgh, on behalf of the Sloan Digital

Sky Survey (SDSS) collaboration, reports the discovery of a type-Ia supernova, ≈ 9 days after maximum light, on a 45-min spectrogram (range 380–980 nm) taken with the 2.5-m telescope (+ multi-fiber spectrograph) at Apache Point Observatory in the course of the SDSS main galaxy redshift survey around Feb. 20.27 UT. The following SDSS position of SN 2004ar is the centroid of the galaxy targeted by the spectroscopic observation (no imaging was obtained, unfortunately, though an SDSS image taken on 2003 Dec. 25 may show the supernova faintly), the supernova estimated as being likely within 1′′.5 of the targeted position: α = 9^{h}59^{m}48^{s}.1, δ = +11^{o}28′25′′ (equinox 2000.0). The redshift of the host galaxy is z = 0.0604. The estimated apparent magnitudes of SN 2004ar are g = 18.6, r = 18.34, and i = 18.6 (based on the spectrophotometry and measured within the 3′′-diameter aperture centered on the galaxy). Following a request from the Central Bureau, P. Holvorcem and M. Schwartz took a 240-s unfiltered (red) CCD exposure with the Tenagra 0.81-m telescope on Mar. 11.25, which shows a point-source object of mag 17.9 (that is not visible on the red Digitized Sky Survey image) located at α = 9^{h}59^{m}48^{s}.17, δ = +11^{o}28′26′′.5, which is 0′′.9 west and 2′′.6 north of a nearby galaxy.

Further to IAUC 8300, H. Pugh, J. Burket, and W. Li report the LOSS discovery of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004as Mar. 11.31 11^{h}25^{m}39.^{s}18 +22^{o}49′49′′.4 17.3 4′′.7 E, 2′′.9 S
2004at Mar. 12.30 10 58 45.18 +59 29 12.0 17.7 12′′.6 W, 0′′.1 S

Additional KAIT magnitudes: SN 2004as, Feb. 23.41 UT, [18.0; Mar. 10.31, 17.5; 12.38, 17.2. SN 2004at in MCG +10-16-37, Feb. 29.35, [18.5, Mar. 13.29, 17.2.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004as, obtained by P. Berlind on Mar. 13.30 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova a few days before maximum light. A narrow Hα emission line superposed on the spectrum of the supernova indicates a recession velocity of 9300 km/s for the host galaxy. Using this value, the supernova expansion velocity is ∼ 14100 km/s for Si II (rest 635.5 nm).

2004 March 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green